Determine the first order partial derivative of the following:
1. F(x, y) = cos(e^t) dt
2.f(x,y,z)=xy^2e^-xz
1
Expert's answer
2021-08-03T15:17:12-0400

1.

Given that

"F(x,y)=\int_y ^xcos(e^t)dt"

Differentiating partially with respect to x we get

"\frac{\delta F}{\delta x}=\frac{\delta}{\delta x}\int_y^xcos(e^t)dt"


"=cos(e^x).\frac{\delta}{\delta x}(x)+cos(e^t).\frac{\delta}{\delta x}(y)................"


By leibritz theorem

"\frac{\delta F}{\delta x}=cos(e^x)"


now, again differenting partially with respect to y we get

"\frac{\delta F}{\delta y}=\frac{\delta}{\delta y}\int_y^xcos(e^t)dt"


"=cos(e^x).\frac{\delta}{\delta y}(x)-cos(e^y).\frac{\delta}{\delta y}(y)"


by Leibritz rule

"\frac{\delta F}{\delta y}=-cos(e^y)"


2.

Given "f(x,y,z)=xy^2e^{-xz}"

The first order derivative

"f(x,y,z)=xy^2e^-{xz}\\F_x=\frac{\delta}{\delta x}(xy^2e^{-xz})\\y^2\frac{\delta}{\delta x}(xe^{-xz})\\y^2[x\frac{\delta}{\delta x}e^{-xz}+e^{-xz}\frac{\delta}{\delta x} x]\\F_x=y^2[xe^{-xz}(-z)+e^{-xz}]"

"F_x(x,y,z)=e^{-xz}y^2(1-xz)"

"F_y=\frac{\delta}{\delta y}(xy^2e^{-xz})=F_y=2xye^{-xz}\space and \space F_z=\frac{\delta}{\delta z}(xy^2e^{-xz})"

"F_z=xy^2\frac{\delta}{\delta z}(e^{-xz})\\xy^2e^{-xz}\frac{\delta}{\delta z}(-xz)\\F_z=-x^2y^2e^{-xz}"